(x-2)^2=49
Find all the real or imaginary solutions to the equation.

Answer by stanbon(75887)

(x-2)^2=49
Find all the real or imaginary solutions to the equation.
---------------
(x-2)^2-49=0
(x-2-7) = 0 or (x-2+7)=0
x=9 or x=-5
===========================
Cheers,
Stan H.